L. Shapakidze

A. Razmadze Mathematical Institute, Georgian Academy of Sciences Tbilisi, Georgia

ON BIFURCATION OF SECONDARY FLOWS OF VISCOUS

INCOMPRESSIBLE FLUID BETWEEN TWO POROUS ROTATING CYLINDERS

In the present report we consider the flow of viscous incompressible liquid between two counter-rotating porous cylinders. The oscillatory modes and bifurcation appearing in the vicinity of the intersection point of neutral curves, corresponding to rotating symmetrical and oscillatory instability is investigated.